Given the function, f(x,y) = x^3 y + 12x^2 - 8y Find all critical points and classify each as a local maximum, local minimum, or neither.


1
Expert's answer
2021-10-12T04:41:08-0400

Given function is "f(x,y)=x^3y+12x^2-8y=0"

Differentiating the function w.r.t. "(x)," we get:

"3x^2y+x^3\frac{dy}{dx}+24x-8\frac{dy}{dx}=0\\
\Rightarrow \frac{dy}{dx}=\frac{24x+3x^2y}{8-x^3}"

For critical points, put "\frac{dy}{dx}=0," we get:

"\frac{24x+3x^2y}{8-x^3}=0\\
\Rightarrow 24x+3x^2y=0\\
\Rightarrow 3x(8+xy)=0"

"\therefore x=0 \ or \ xy=-8"


Case 1: If "x=0" , then:

"f(x,y)=-8y=0\\
\Rightarrow y=0"

Case 2: If "xy=-8" , then:

"f(x,y)=-8x^2+12x^2-8y=0\\
\Rightarrow 4x^2-8y=0\\
\Rightarrow \frac{x^2}{2}=y"

So, the given function can be written as:

"x^3(\frac{x^2}{2})+12x^2-4x^2=0\\
\Rightarrow \frac{x^5}{2}+8x^2=0\\
\Rightarrow x^2(\frac{x^3}{2}+8)=0\\
\Rightarrow x=0 \ or\ x^3=-16\\
\Rightarrow x=0 \ or\ x=-(16)^{\frac{1}{3}}"

"\therefore y=0 \ or\ y=\frac{(16)^{\frac{2}{3}}}{2}"

So, from both cases, we obtain the same point: "(0,0)" and "(-(16)^{\frac{1}{3}}, \frac{(16)^{\frac{2}{3}}}{2})"


Now, calculating double derivative, we get:

"6xy+3x^2\frac{dy}{dx}+3x^2\frac{dy}{dx}+x^3\frac{d^2y}{dx^2}+24-8\frac{d^2y}{dx^2}=0\\
\Rightarrow \frac{d^2y}{dx^2}=\frac{6xy+6x^2\frac{dy}{dx}+24}{8-x^3}"

Now, calculating value of double derivative at "(0,0)" , we get:

"(\frac{d^2y}{dx^2})_{(0,0)}=\frac{24}{8}=3" which is positive. Hence, the point is local minimum.


Now, calculating value of double derivative at "(-(16)^{\frac{1}{3}}, \frac{(16)^{\frac{2}{3}}}{2})", we get:

"(\frac{d^2y}{dx^2})_{(-(16)^{\frac{1}{3}}, \frac{(16)^{\frac{2}{3}}}{2})}=\frac{-48+0+24}{8+16}=-1" which is negative. Hence, the point is local maximum.